# Constant-Memory Strategies in Stochastic Games: Best Responses and Equilibria

### v1

@@ 4.2 Computing Best Responses to Mixed Strategies is Intractable
Solving CMDP reduces to computing the best response for agent $i$ to a mixed stationary strategy profile $(\Pi^{0+}_{-i},\vec{p})$ of the opponents.
@@
We prove the above theorem by constructing a reduction for every CMDP instance to a BestResponse instance. Given a CMDP $\langle\mathcal{C},\mathcal{S},\mathcal{A},f_{T},f_{R},\gamma\rangle$, where
  1. 1.
    $\mathcal{C}$ is a finite set of unobservable contexts, one of which will be selected at the beginning of an episode.
  2. 2.
    $f_{T}$ and $f_{R}$ take in a context $c\in\mathcal{C}$ and output a transition function $T^{c}:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ as well as a reward function $R^{c}:\mathcal{S}\times{A}\mapsto\mathbbm{R}$, respectively. $\langle\mathcal{S},\mathcal{A},T^{c},R^{c},\gamma\rangle$ will then constitutes an MDP.
one can construct a stochastic game with two players $\langle\{1,2\},\mathcal{S},\{\mathcal{A},\mathcal{A^{\prime}}\},T_{0},\{R_{1},R_{2}\}\rangle$, where agent 1 with action set $\mathcal{A}$ is playing against agent 2 (as a context controller/switcher) with a set of stationary strategies $\{\pi_{c}:\mathcal{S}\mapsto\Delta(\mathcal{A}^{\prime})\}_{c\in\mathcal{C}}$. That is, there exists a $T_{0}$ and $\{\pi_{c}\}_{c\in\mathcal{C}}$ such that the following system of equations holds simultaneously.
$\forall c\in\mathcal{C},\ T^{c}(S^{\prime}|S,a)=\sum_{a^{\prime}}T_{0}(S^{\prime}|S,a,a^{\prime})\pi_{c}(a^{\prime}|S)$ (14)
One can see that the above equation operates independently for each $(S,a)$ pair, but should hold simultaneously for all $c\in\mathcal{C}$ while fixing a pair of $(S,a)$.
For each $(S,a)$, in a matrix notation, Equation (14) can be written as
$M_{C}=M_{\Pi}\cdot M_{T}$
where $M_{C}[c,S^{\prime}]=T^{c}(S^{\prime}|S,a),M_{\Pi}[c,a^{\prime}]=\pi_{c}(a^{\prime}|S),M_{T}[a^{\prime},S^{\prime}]=T_{0}(S^{\prime}|S,a,a^{\prime})$, and therefore, $M_{C}\in\mathbbm{R}^{C\times S},M_{\Pi}\in\mathbbm{R}^{C\times A^{\prime}},M_{T}\in\mathbbm{R}^{A^{\prime}\times S}$. That is, every row $r_{c}\in\mathbbm{R}^{1\times S}$ of $M_{C}$ is a linear combination of all rows in $M_{T}$, and the linear weights is given as the $c$-th row $\in\mathbbm{R}^{1\times A^{\prime}}$ of $M_{\Pi}$. Then, a trivial way to find such $M_{\Pi}$ and $M_{T}$ is as follows,
  1. 1.
    If $|C|\leq|S|$, then one can make $|A^{\prime}|=|C|$, and $M_{\Pi}=\mathbf{I}_{C\times C}$ and $M_{T}=M_{C}$. That is, the policy of the context switcher is to simply select a context in a deterministic way.
  2. 2.
    If $|C|>|S|$, to avoid $M_{\Pi}$ and $M_{T}$ being unboundedly large, a straightforward way is to make $M_{T}=\mathbf{I}_{S\times S}$, i.e. $|A^{\prime}|=|S|$, as the dimension of the row space of $M_{C}$ is at most $|S|$, and therefore, a set of orthonormal basis vectors (i.e., each row of $M_{T}$) can already span the whole row space of $M_{C}$.
One may note that even if in the sub-case when $|C|\leq|S|$, having $|A^{\prime}|=|C|$ may be redundant. A natural question is to find such $M_{\Pi}$ and $M_{T}$ of minimum sizes, i.e. with smallest $|A^{\prime}|$. This further reduces to finding a minimum set of $|S|$-dimensional vectors whose linear combination can represent all the row vectors in $M_{C}$. We now describe a procedure that iteratively construct such $M_{\Pi}$ and $M_{T}$, formally given in Algorithm 1. The idea is quite clean and elegant: start with any arbitrary row in $M_{C}$ as the first basis and an ordering of the rest rows, project the $i$-th subsequent row onto all previous $(i-1)$ basis vectors, and treat the orthogonal residual as the $i$-th basis vector if it is non-zero. One may note that this procedure resembles the Gram-Schmidt Orthogonalization [14], which can be done in strongly-polynomial time. The only difference is, the standard Gram-Schmidt Orthogonalization starts with a set of vectors that are already linearly independent but might not be orthogonal to each other, while here we start with a set of vectors that are possibly linearly dependent and the goal is to find the minimum set of basis vectors.
Algorithm 1 Find the minimum $M_{\Pi}$ and $M_{T}$
1:  Input: $M_{C}$ with $|C|\leq|S|$
2:  Output: $M_{\Pi}$ and $M_{T}$ of the minimum size
3:  Initialize: $M_{T}$ as an empty matrix
4:  $M_{T}.append\_row(M_{C}[1])$
5:  for $i=2$ to $|C|$ do
6:     $new\leftarrow M_{C}[i]-\sum_{k=1}^{i-1}\frac{\langle M_{T}[k],M_{C}[i]\rangle}{\langle M_{T}[k],M_{T}[k]\rangle}M_{T}[k]$
7:     if $new\neq\vec{0}$ then
8:        $M_{T}.append\_row(new)$
9:     end if
10:  end for
11:  $M_{T}\leftarrow normalize\_each\_row(M_{T})$
12:  $M_{\Pi}\leftarrow M_{C}\cdot transpose(M_{T})$
13:  return  $M_{\Pi}$, $M_{T}$
@@
It has been proved that in infinitely repeated games, there exists a Turing machine strategy such that no Turing machine can implement its best response [27, 35]. Based on the widely accepted belief that CMDP is in the same complexity class as POMDP, Theorem 9 further implies that to render the best response computationally undecidable, it is probably sufficient for the opponent to mix over just a few constant-memory strategies, rather than implementing a complex strategy with perfect memorization via a computer program.
Another remark is, as indicated by Equation (13), if in each turn the opponent is allowed to switch to a different strategy independently of previous actions, then solving such a belief-induced MDP is sufficient for computing the best response, which is definitely decidable and even has constant-memory strategies as optimal policies.
5 Revisiting Single-Agent Decision Processes: A Generative Framework
Building on the theoretical insights from the previous discussion, we now present a generative framework for creating single-agent learning environments. RL is notorious for its poor generalization ability even within the same domain, as an RL model is typically trained for a particular environment simulated by an underlying fixed MDP-like process. Since generalizability is crucial in many real-world applications, especially when an RL algorithm is deployed, the community has investigated some similar research ideas under various names, e.g., multi-task RL [50, 53, 52], contextual RL [23, 11], epistemic POMDPs [20], and zero-shot generalization [26].
Therefore, a framework that can automatically generate environment instances with some shared features is of great significance, since it can be utilized for both training and evaluating the generalizability of a single-agent RL algorithm. There are a few exemplars, e.g. using procedural content generation to make level- or seed-based video games (ProcGen [16]), or simulating existing RL environments with user-specified contextual parameters (CARL [^1] https://github.com/automl/CARL [11]). As a supplement, here we also provide a framework for generating for generating instances of single-agent RL environments. More precisely, this can be done by specifying opponents with controlled memory (hence controlled bounded rationality) under a multi-agent skeleton simulator. The induced decision-making process for the remaining agent will then be a highly configurable learning environment. A key advancement of our generative framework is that the induced single-agent environments can be arbitrarily hard to learn, while the aforementioned existing benchmarks in the literature only diversify tasks with comparable levels of difficulty.
As we addressed from the theoretical side, optimally solving the induced single-agent decision problem involves best responding to the underlying hidden opponents, in other words, the reward/context controller. Once we have an underlying skeleton simulator as a stochastic game (e.g., Figure 1), and specify a context/reward controller as an opponent, we can generate a single-agent learning environment that is arbitrary hard to learn through the following procedures, in ascending order of difficulty:
  1. 1.
    If the opponent plays a single stationary strategy, this will lead to a standard RL environment.
  2. 2.
    If the opponent plays a single behavioral $K$-memory strategy, this will result in an RL environment that requires the learner to at least stack the past $K$ observations, e.g., Figure 2.
  3. 3.
    If the opponent plays a mixed $K$-memory strategy, this will lead to an RL environment that requires the learner to leverage a Bayesian adaptive policy, e.g., Figure 3.
1class StochasticGame:
2 def __init__(self, ...):
3 # instantiate a stochastic game
4 pass
5
6 def reset(self):
7 # initialize a new match, and return the initial state
8 return obs_n, info_n
9
10 def step(self, action_n):
11 # proceed the game to the next state
12 return obs_n, reward_n, term_n, trunc_n, info_n
Figure 1: Standard Pythonic interfaces of a multi-agent simulator as a stochastic game.
1class KMemoryOpponentWrapper:
2 def __init__(self, stochastic_game, agent_i, pi_others, K):
3 # take in a stochastic game as the underlying skeleton
4 self.sg = stochastic_game
5 self.agent_i = agent_i
6 self.pi_others = pi_others
7 self.K = K
8
9 def reset(self):
10 # initialize a new match, and return the initial state
11 obs_n, info_n = self.sg.reset()
12 self.hist_K = [obs_n]
13 return obs_n[self.agent_i], info_n[self.agent_i]
14
15 def step(self, action_i):
16 # proceed the game to the next state
17 action_others = self.pi_others(self.hist_K)
18 obs_n, reward_n, term_n, trunc_n, info_n =\
19 self.sg.step((action_i, action_others))
20 self.hist_K = (self.hist_K + [obs_n])[-self.K:]
21 return (obs_n[self.agent_i], reward_n[self.agent_i],
22 term_n[self.agent_i], trunc_n[self.agent_i],
23 info_n[self.agent_i])
Figure 2: Single-agent learning environments induced by a behavioral $K$-memory opponent strategy.
1class MixedKMemoryOpponentWrapper(KMemoryOpponentWrapper):
2 def __init__(self, stochastic_game, agent_i,
3 pi_others_set, probs, K):
4 # take in a stochastic game as the underlying skeleton
5 super().__init__(stochastic_game, agent_i, None, K)
6
7 # now a set of strategies and the corresponding probs
8 self.pi_others_set = pi_others_set
9 self.probs = probs
10
11 def reset(self):
12 # initialize a new match, and return the initial state
13 obs, info = super().reset()
14 self.pi_others =\
15 random.choice(self.pi_others_set, p=self.probs)
16 return obs, info
Figure 3: Single-agent learning environments induced by a mixed $K$-memory opponent strategy.
6 Conclusion and Future Work
In this work, we comprehensively investigate the concept of $K$-memory strategies. We first establish the best responses and Nash equilibria for behavioral $K$-memory strategies, followed by a discussion on the computational hardness of best responding to mixed $K$-memory strategies. Those theoretic insights later empower a generative framework for studying generalizability of single-agent RL algorithms.
A few research problems still remain open, e.g.,
  1. 1.
    How to efficiently synthesize a truly $K$-memory strategy (ensuring it does not accidentally reduce to a $(K-1)$-memory one), and how to compactly represent it (instead of enumerating all the entries)? Chances are that these two lie in two ends of a spectrum.
  2. 2.
    What can be the minimum effort to make a constant-memory strategy even more powerful? For example, in the Iterated Prisoner’s Dilemma, if both agents adopt a cyclic strategy (constant-memory) together with a Grim-Trigger (definitely not constant-memory but way easy to simulate using a binary indicator), then any payoff profile (under the average-payoff criterion though) can be realized as the outcome of an equilibrium [44].
In summary, we hope this work provide a concise way to characterize bounded rationality in agents that learn and interact, benefiting various real-world applications, such as human-robot cohabitation, strategic LLM-agents, and other domains in cognitive science. We will report more comprehensive experimental results in the near future.

### v2

@@ D.1. Generalized (One-Shot) Traveler’s Dilemma
$TD(2,2,[2..100])$ has a generalized ordinal potential.
@@
We prove it by construction. We will construct such a $P$ divided by cases.
  1. (1)
    If $a_{1}=a_{2}=\alpha$, then $P(\alpha,\alpha)=2\times(101-\alpha)-1$
  2. (2)
    If $a_{1}=a_{2}-1=\alpha-1$, then $P(\alpha-1,\alpha)=2\times(101-\alpha)$. Same for $P(\alpha,\alpha-1)$.
  3. (3)
    If $a_{1}=a_{2}-1-x=\alpha-1-x$, then $P(\alpha-1-x,\alpha)=2\times(101-\alpha)-x$. Also, same for $P(\alpha,\alpha-1-x)$.
Figure 5. A general ordinal potential of $TD(2,2,[2..100])$.
One can then examine the definition. W.l.o.g., we might as well consider fixing $a_{2}$, for two possible bids of agent 1, $a_{1}$ and $a_{1}^{\prime}$,
  1. (1)
    If $a_{1}>a_{2}+1$ and $a_{1}^{\prime}>a_{2}+1$, $u_{1}(a_{1},a_{2})=u_{1}(a_{1}^{\prime},a_{2})=a_{2}-2$. The premise of the implication is not satisfied, hence the implication is vacuously true.
  2. (2)
    If $a_{1}=a_{1}^{\prime}$, the definition is also vacuously true.
  3. (3)
    If $a_{1}=a_{2}-1$, then $u_{1}(a_{1},a_{2})=a_{2}-1+2=a_{2}+1$, and $P(a_{1},a_{2})=2\times(101-a_{2})$, by the following sub-cases,
    1. (a)
      If $a_{1}^{\prime}>a_{2}+1$, then
      $\begin{cases}u_{1}(a_{1}^{\prime},a_{2})=a_{2}-2\\ P(a_{1}^{\prime},a_{2})=2\times(101-a_{1}^{\prime})-(a_{1}^{\prime}-1-a_{2})=202-3a_{1}^{\prime}+a_{2}+1<202-2a_{2}-2\end{cases}$
    2. (b)
      If $a_{1}^{\prime}=a_{2}+1$, then
      $\begin{cases}u_{1}(a_{1}^{\prime},a_{2})=a_{2}-2\\ P(a_{1}^{\prime},a_{2})=2\times(101-a_{1}^{\prime})=202-2a_{2}-2\end{cases}$
    3. (c)
      If $a_{1}^{\prime}=a_{2}$, then
      $\begin{cases}u_{1}(a_{1}^{\prime},a_{2})=a_{2}\\ P(a_{1}^{\prime},a_{2})=2\times(101-a_{2})-1=202-2a_{2}-1\end{cases}$
    4. (d)
      If $a_{1}^{\prime}=a_{2}-1$, then $a_{1}=a_{1}^{\prime}$, and the definition is vacuously true.
    5. (e)
      If $a_{1}^{\prime}<a_{2}-1$, then
      $\begin{cases}u_{1}(a_{1}^{\prime},a_{2})=a_{1}^{\prime}+2<a_{2}+1\\ P(a_{1}^{\prime},a_{2})=2\times(101-a_{2})-(a_{2}-1-a_{1}^{\prime})=202-3a_{2}+a_{1}^{\prime}+1<202-2a_{2}\end{cases}$
    All the above sub-cases satisfy $P(a_{1},a_{2})>P(a_{1}^{\prime},a_{2})$ whenever $u_{1}(a_{1},a_{2})>u_{1}(a_{1}^{\prime},a_{2})$.
The proof of Theorem 16 provides one with an illustrative example where a generalized ordinal potential function $P$ can be directly devised along with its rough structure. Further, we show in general the multi-player TD game has a generalized ordinal potential, however, by a powerful but rather intuitive lemma so as to make our lives much easier, instead of explicitly coming up with a desired function $P$.
Definition 0 (Improvement path).
An improvement path with respect to a bidding profile a\@vec{a} is a maximal sequence of profitable deviations starting with a\@vec{a}.
Lemma 0 (Finite improvement property).
A game is said to have the finite improvement property (FIP) if every improvement path is finite. Every finite game has a generalized ordinal potential iff it has the FIP.
@@
Any $TD(N,k,A)$ has a generalized ordinal potential.
Proof.
Starting from any bidding profile, a profitable deviation is for a bidder that is not currently the lowest bidder (including tied bids) to bid less than or equal to the currently lowest one. The improvement path cannot be any longer if every one reaches the lowest possible bid. ∎
D.2. The Iterated Traveler’s Dilemma
An iterated traveler’s dilemma (ITD) consists of a Markov chain of TDs. More specifically, an ITD is a 5-tuple $\langle N,k,[\underline{A}..\overline{A}],p,\gamma\rangle$, given as follows,
  1. (1)
    For all $A\subseteq[\underline{A}..\overline{A}]$, $\langle N,k,A\rangle$ is a valid TD defined as previously, i.e. $TD(N,k,A)$.
  2. (2)
    Given the current dilemma $TD(N,k,[A_{1}..A_{2}])$, and a bidding profile $\@vec{a}\in[A_{1}..A_{2}]^{N}$, the successor dilemma can be changed to $TD(N,k,[op_{1}(\@vec{a})..op_{2}(\@vec{a})])$ w.p. $p$, or stay the same otherwise. In particular, $op_{1}$ and $op_{2}$ are the two operators that aggregates the last bidding profiles. For example, $op_{1}(\cdot)$ can a constant function that always outputs $\underline{A}$, and $op_{2}(\cdot)=\max(\cdot)$, implying an ITD whose available bidding space may shrink. Formally, a transition function is defined as
    $T\Big(S^{\prime}\Big|S=TD(N,k,[A_{1}..A_{2}]),\@vec{a}\Big)=\begin{cases}p,&\text{ if }S^{\prime}=TD(N,k,[op_{1}(\@vec{a})..op_{2}(\@vec{a})])\\ 1-p,&\text{ if }S^{\prime}=TD(N,k,[A_{1}..A_{2}])\\ 0,&\text{ otherwise}\end{cases}$
  3. (3)
    A reward function that assigns each agent an immediate signal is defined as
    $R_{i}\Big(TD(N,k,A),\@vec{a}\Big)=\begin{cases}u_{i}(\@vec{a};TD(N,k,A)),&\text{ if }a_{1}=a_{2}=\cdots=a_{n}\\ 0,&\text{ otherwise}\end{cases}$
  4. (4)
    The game ends immediately after when $a_{1}=a_{2}=\cdots=a_{n}$, and the rewards of this final round will also be collected. The overall objective for agent $i$ is to maximize the discounted accumulated rewards $\sum_{t=0}^{T-1}\gamma^{t}R_{i}^{t}$, where $T$ is the number of transitions in this episode.
The game will end according to some rules, e.g., when the number of rounds exceeds a specified number (as in the main body of this paper), or when all agents bid the same (a way harder version).
D.3. Sample Curves of the Training Phase
With some main results summarized in Figure 2, we here also provide the detailed training curve, as shown in Figure 6.
Figure 6. Training process for the Iterated Traveler’s Dilemma. Each sub-figure denotes the best response (using various memory lengths) against a particular length of the memory used by the opponent.
Appendix E Details for the Pursuit Domain
Figure 7. An illustration of the Pursuit domain, where red dots are pursuers and blue dots are evaders. The orange squared centered at each red dot is the observed local area of that agent.
E.1. Detailed Experimental Settings
E.1.1. Environment Setup
The Pursuit testbed Gupta et al. (2017), illustrated in Figure 7, allows the users to specify a few parameters, in order to deliver a customized environment. We have involved four configurations, as listed in Table 2. Specifically, we use the first three configurations to benchmark different RL algorithms in the next subsection, while the fourth configuration is used in the experiment in the main text.
Config max_cycles width height #evaders #pursuers obs_range tag_reward catch_reward urgency_reward
HighCatch 300 10 10 10 8 7 0.1 5 -0.1
HighTag 2 0.1 -0.1
SameTagCatch 1 1 -0.1
HighCatchLarge 16 16 20 8 7 0.1 3 -0.1
Table 2. The detailed parameters of each configurations involved
E.1.2. Hardware
We use Linux Servers with NVIDIA GeForce RTX 3090 GPUs.
E.1.3. Network Architecture
All RL policies are equipped with a convolutional preprocessing network. For this preprocessing network, we use three sequential Convolution layers, namely
  1. (1)
    Conv2d(input_channels=3, output_channels=32, kernel_size=4, stride=1, padding=0),
  2. (2)
    Conv2d(input_channels=32, output_channels=64, kernel_size=2, stride=1, padding=0),
  3. (3)
    Conv2d(input_channels=64, output_channels=64, kernel_size=2, stride=1, padding=0),
with a ReLU activation followed after each layer. Finally all the features are flattened and projected into a 512-dimensional vector. For main body of the policy/value network, we use three-layer MLPs of the hidden dimension [512, 256, 256].
E.1.4. RL Algorithm Parameters
We mainly adopt the implementation by Raffin et al. (2021). For learning to find NEs under multi-agent settings, we equip each agent with the same network and RL algorithm, and make them learn independently. Specially, for the detail algorithmic parameters,
  1. (1)
    In DQN, we have batch_size=256, exploration_fraction=0.2;
  2. (2)
    In PPO, we have batch_size=256, making the policy/value network of the same architecture but updated independently;
  3. (3)
    In A2C, we have ent_coef=0.01, vf_coef=0.5, n_steps=400, and same network setup with PPO.
Other parameters that are not mentioned are set to be their default values.
E.2. Benchmarking Different Deep RL Algorithms
Figure 8. Benchmarking results using different algorithms and memory of varying lengths for the Pursuit domain instantiated under three configurations. From left to right, each figure represents the experimental results for the configuration HighCatch, SameTagCatch, and HighTag, respectively.
Additional to the results presented in the main text that are obtained using DQN. We benchmark all three algorithms, including the other two, namely PPO and A2C, with the results summarized in Figure 8. Among these three, DQN performs always the most effectively:
  1. (1)
    For HighCatch, DQN can always effectively make the hunting process shorter. A longer-memory setting leads to a shorter period of hunting. In contrast, PPO and A2C do not result in successful cooperative hunting strategies.
  2. (2)
    For SameTagCatch, the way we set-up the rewards shall lead to a solution where each agents is supposed to first tag evaders on its own, but ends up catching and removing them under cooperation with the pursuers. Therefore, DQN agents equipped with long memories tend to hunt those evaders more “slowly”, leaving more time for themselves to tag the evaders to obtain sufficient rewards before the game terminates. In contrast, we found that although the episode lengths of PPO and A2C remain high, they are not learning to tag agents, indicated by their low returns during training.
  3. (3)
    For HighTag, agents shall figure out that cooperatively catching evaders is not a desired strategy; rather, independently tagging the evaders without removing them from the game is supposed to be the best strategy. Thus, in most cases, the agents operates for the full episode. Except for one case of DQN with 8-memory, it seems that the agent is still confused about whether to adopt a tagging-without-hunting strategy.